Solve using properties of logarithms:
log2 (x+1)- log2 x = log2 5
note- the numbers directly following "log" are suppossed to be the spaces. I do not know how to type them to be small and at the bottom

Answer by venugopalramana(3286)

Solve using properties of logarithms:
log2 (x+1)- log2 x = log2 5
note- the numbers directly following "log" are suppossed to be the spaces. I do not know how to type them to be small and at the bottom
YOU MEAN ALL LOGS ARE TO BASE 2....IF ALL LOGS ARE TO SAME BASE ITDOES NOT MATTER IF WE OMIT WRITING THAT BASE...SINCE LOG X TO BASE TO LOG 2 = LOG X/LOG 2...LIKE THAT EVERY TERM HAS LOG 2 IN THE DENOMINATOR ON BOTH SIDES WHICH WE CAN CANCEL OUT. SO WE GET
LOG(X+1)-LOG X=LOG 5
USING LOG A - LOG B = LOG (A/B) ..WE GET...
LOG(X+1)/X = LOG 5 ...TAKING ANTILOGS ,WE GET ..
(X+1)/X=5
5X=X+1
5X-X=1
4X=1
X=1/4